OpenStudy (anonymous):

What is the vertex of the graph of y = 2(x – 3)2 + 4? (4, 3) (4, –3) (3, 4) (–3, 4)

OpenStudy (mertsj):

(3,4)

jimthompson5910 (jim_thompson5910):

The vertex of y = a(x-h)^2+k is (h,k) In the case of y = 2(x – 3)2 + 4, h = 3 and k = 4 So the vertex is (3, 4)